OpenStudy (anonymous):

Graph the equation -3x + 4y = 12 Please help me solve and show steps.

OpenStudy (anonymous):

find y

OpenStudy (anonymous):

First let's pit into straight line form: So: 4y=12+3x y=3+2/4x

OpenStudy (anonymous):

y=3+3/4x

OpenStudy (anonymous):

Then look for the intercepts.

OpenStudy (anonymous):

To find the x intercept set y to zero and solve. To find the y intercept set x to zero and solve

OpenStudy (anonymous):

Thanks I'll give it a try!

OpenStudy (anonymous):



OpenStudy (anonymous):

you have slope there which is 3/4..start y =3 and go up 3 and over 4

OpenStudy (anonymous):

You can tell which way the line will go based on whether the gradient - m - is positive or negative.

OpenStudy (anonymous):

Where does the 2 come from?

OpenStudy (anonymous):